# Gender Bias in Perception of Human Managers Extends to AI Managers

### v2

@@ Abstract
As AI becomes more embedded in workplaces, it is shifting from a tool for efficiency to an active force in organizational decision-making. Whether due to anthropomorphism or intentional design choices, people often assign human-like qualities, including gender, to AI systems. However, how AI managers are perceived in comparison to human managers and how gender influences these perceptions remains uncertain. To investigate this, we conducted randomized controlled trials (RCTs) where teams of three participants worked together under a randomly assigned manager. The manager was either a human or an AI and was presented as male, female, or gender-unspecified. The manager’s role was to select the best-performing team member for an additional award. Our findings reveal that while participants initially showed no strong preference based on manager type or gender, their perceptions changed significantly after experiencing the award process. As expected, those who received awards rated their managers as more trustworthy, competent, fair, and were more willing to work with similar managers in the future, while those who were not selected viewed them less favorably. However, male managers, whether human or AI, were more positively received by awarded participants, whereas female managers, especially female AI managers, faced greater skepticism and negative judgments when they did not give awards. These results suggest that gender bias in leadership extends beyond human managers to include AI-driven decision-makers. As AI assumes more managerial responsibilities, understanding and addressing these biases will be crucial for designing fair and effective AI management systems.
@@ 1 Introduction
The increasing integration of artificial intelligence (AI) in workplaces is transforming its function from merely enhancing efficiency to actively shaping organizational decision-making, sparking discussions on the evolving dynamics between humans and machines [1]. AI algorithms are increasingly taking on roles traditionally performed by humans, such as assisting with decisions related to promotion and hiring [2, 3]. While human managers often retain final authority, some companies are experimenting with AI in leadership-adjacent roles, such as managing repetitive tasks or optimizing team operations [4]. This growing influence of AI raises questions about workplace dynamics, particularly how AI decision-makers are perceived relative to human managers. Issues such as fairness, biases, organizational justice [5], trust, morale, and willingness to collaborate within teams are becoming increasingly important as AI reshapes the future of leadership and management.
@@
Even when AI lacks explicit gender markers, individuals tend to subconsciously assign gender based on factors like language style, voice, and the perceived roles of AI. Research shows that minimal gender cues, such as vocal cues embedded in a machine, can evoke gender-based stereotypical responses [36]. A study on ChatGPT shows it is more likely to be perceived as male than female due to its competence-oriented attributes, such as providing information or summarizing text [25]. AI systems designed for caregiving or support, such as digital assistants like Siri and Alexa, tend to be feminized through their anthropomorphization [37], reflecting societal norms that associate caregiving and service roles with women [38]. The subconscious gendering of AI can influence fairness perceptions and team dynamics, making it important to understand these biases to promote positive interactions in human-AI collaborations.
@@
The interplay between manager type (human or AI) and perceived gender, in shaping perceptions of managers, particularly in competitive scenarios where only a subset of individuals are awarded, remains underexplored. While gender bias is a central concern of this research, it is part of a broader investigation into how multiple interacting factors influence managerial perceptions. This study aims to examine how the combined effects of manager type, manager gender, self-perceived contribution, and award status influence the perception of managers’ trustworthiness, competence, fairness, and willingness to work with similar managers in future tasks. Trustworthiness relates to reliability and ethical standing; competence underpins a manager’s authority and efficacy; fairness is crucial for morale and acceptance; and willingness to work in future endeavors reflects the sustainability of the collaboration. These dimensions are essential for fostering positive and effective human-AI team interactions.
To explore these dynamics, we conducted online experiments where participants worked in teams under managerial scenarios varying by manager type and gender. This study provides insights into the nuanced ways that both AI and gendered presentations impact human acceptance of AI managers, ultimately contributing to the growing understanding of effective AI integration in workplaces. Based on prior literature and theoretical frameworks, we propose the following hypotheses:
  • H1: Participants who receive an award will rate their manager as more trustworthy, competent, fair, and will express greater willingness to work with a similar manager in the future than participants who do not receive an award.
@@ 2 Materials and Methods
The survey questions used in the pre-treatment and post-treatment surveys, along with screenshots from the experiment, are provided in the Supporting Information. This research complies with University College Dublin (UCD) Human Research Ethics Regulations and Human Research Ethics Committee (HREC) guidelines for research involving human participants. The research study protocol has been approved by UCD Human Research Ethics (HS-LR-24-10-Cui-Yasseri). Informed consent was obtained from all human participants prior to the experiment.
@@ 2.2 Design and Measures
Figure 1 illustrates the timeline view of the experimental methodology. This study employed a mixed experimental design to examine how participants’ perceptions of trustworthiness, competence, fairness, and willingness to work with a manager were influenced by primary factors: manager type, manager gender, award outcome, and self-perceived contribution. Participants were randomly assigned to three-person groups, with each group placed in a different manager condition. In these conditions, participants viewed scenarios in which either a human or an AI manager made a selection decision. Within each manager type, the gender of the manager was also manipulated, with the manager identified as either female, male, or gender unspecified to explore potential biases or differential expectations based on perceived gender.
@@
The dependent variables in this study were the participants’ perceptions of trustworthiness, competence, fairness of their assigned manager, and the willingness to collaborate with a similar manager in future tasks. Each of these perceptions was measured using a Likert-type scale from 0 to 10, with items such as “How fair do you think the manager was in making the award decision in the game?”. Take fairness as an example; we employed the following linear regression model to examine the factors influencing post-treatment fairness perceptions:
@@ 2.3 Procedure
Participant onboarding and pre-treatment survey. The experiment took place entirely online and was designed to simulate a decision-making scenario involving teamwork and managerial evaluation. Participants recruited via Prolific began by clicking a provided link to enter the Empirica platform, where the experiment was hosted. Upon joining, participants first read an information sheet outlining the study’s goals and procedures, followed by providing their consent. They then completed a brief survey collecting demographic information, such as age, gender, education level, and their initial perceptions of trustworthiness, competence, fairness, and willingness to work with both AI and human managers, as well as male and female managers. This initial survey helped establish baseline perspectives toward different manager types.
@@
Individual problem-solving task. After completing the preliminary survey, participants began the first round of the game. A countdown timer appeared at the top center of their screen, along with their randomly assigned ID. They were required to individually complete a visual intelligence task (“Find the Robber” ): read the on-screen description, examine the pictures (SI Fig. S1) based on visual cues, and identify the robber from five options. Participants selected the suspect they believed to be the robber using a slider bar within a 2-minute time limit; otherwise, a default value was automatically recorded.
@@
Manager assignment. After completing the task, participants were informed they would join a team with two teammates for the next round (Figs S2–S5). Each team was randomly assigned a manager, either a human or an AI, who would evaluate the team’s performance. The manager’s gender was also manipulated within each manager type, with the manager identified as male, female, or gender unspecified. The manager would ultimately select the “best player” in the team to receive an extra cash award of £0.5.
@@
Team-based task and collaboration. In the second round, participants collaborated in a team-based setting through an online chat interface to complete the same task as in the first round (SI Fig. S6). They could view their own and their teammates’ responses from the previous round on the slider bars, providing context for discussion. Through the chat, participants worked collectively to arrive at a final team response, simulating real-world team dynamics. All teams completed identical tasks, ensuring a consistent basis for comparing perceptions of the managers across groups.
While the Robber task is relatively simple and does not require advanced math or logic, it engages key cognitive and social processes: perceptual discrimination, judgment under uncertainty (no obvious or guaranteed correct answer), time pressure, and social inference (how participants justify their reasoning and assess others’ input). Despite its simplicity compared to real-world workplace tasks, the Robber task provides a controlled and interpretable model for studying reactions to award decisions in ambiguous, team-based settings where credit attribution is inherently uncertain.
@@
Manager decision and exit survey. After a waiting time of 20 seconds, participants received the announcements of the “best player” selected by the manager (SI Figs. S8–S9). Finally, participants completed an exit survey that included the study’s primary measures: perceptions of trustworthiness, competence, and fairness regarding the manager’s decision, as well as their willingness to work with a similar manager in the future.
Debriefing and payment. After completing the survey, participants were informed during the debriefing that there was no actual manager involved in the experiment and that the selection was random. The entire experiment took approximately 12 minutes, providing an experience of teamwork, evaluation, and managerial decision-making. Participants were compensated at a rate of £9 per hour, prorated for the duration of the study. The chosen “best player” received an additional £0.5 bonus.
@@ 3.1 Comparison of pre- and post-treatment perceptions
Figure 2: Comparison of the mean perceived trustworthiness of different groups by manager type and gender. The upper row represents results for female participants, and the lower row represents results for male participants. (A) and (D) show the mean pre-treatment perceived trustworthiness. (B) and (E) depict the mean change in post-treatment perceived trustworthiness for awarded participants, while (C) and (F) illustrate the change for non-awarded participants. Error bars represent standard error of the mean, $p$-values are from a one-way ANOVA.
Figure 3: Comparison of the mean perceived competence of different groups by manager type and gender. The upper row represents results for female participants, and the lower row represents results for male participants. (A) and (D) show the mean pre-treatment perceived competence. (B) and (E) depict the mean change in post-treatment perceived competence for awarded participants, while (C) and (F) illustrate the change for non-awarded participants. Error bars represent standard error of the mean, $p$-values are from a one-way ANOVA.
Figure 4: Comparison of the mean perceived fairness of different groups by manager type and gender. The upper row represents results for female participants, and the lower row represents results for male participants. (A) and (D) show the mean pre-treatment perceived fairness. (B) and (E) depict the mean change in post-treatment perceived fairness for awarded participants, while (C) and (F) illustrate the change for non-awarded participants. Error bars represent standard error of the mean, $p$-values are from a one-way ANOVA.
Figure 5: Comparison of the mean expressed willingness to work with a similar manager in the future, by manager type and gender. The upper row represents results for female participants, and the lower row represents results for male participants. (A) and (D) show the mean pre-treatment perceived willingness. (B) and (E) depict the mean change in post-treatment perceived willingness for awarded participants, while (C) and (F) illustrate the change for non-awarded participants. Error bars represent standard error of the mean, $p$-values are from a one-way ANOVA.
In the pre-treatment survey, the mean of participants’ perceptions across manager types and genders was around 7, with slight variations (Fig. 25 A, D). Participants generally rated human managers more favorably than AI managers. Female participants gave slightly higher ratings to female managers, and male participants gave slightly higher ratings to male managers, though the overall gender differences were relatively small. The perceptions of managerial decisions show a weak positive correlation between human and AI managers but a strong positive correlation between female and male managers (See SI Fig. S11). This suggests that managerial perceptions in the pre-treatment phase may be more influenced by manager type than by manager gender.
@@
Conversely, not receiving an award (Fig. 25 C, F) generally reduced perceptions of managers, with the steepest declines observed for female-AI managers across both female and male participants, and a statistically significant effect among male participants on all dimensions. The dynamic shifts observed in evaluations before and after the treatment emphasize the influence of manager type, gender, and award outcomes.
@@ 3.2 Effects of demographics and pre-treatment perceptions
Figure 7: Coefficient estimates ($\beta$) for post-treatment perceived competence $Competence_{post}$ among female and male participants, based on pre-treatment perceptions, demographic factors, and interactions. Interaction terms capture the influence of manager type, manager gender, and participants’ self-perceived contributions on perceived competence. Statistical significance is denoted by $p<0.05$. Error bars represent standard errors. Dashed lines visually separate pre-treatment factors, non-awarded AI categories, non-awarded human categories, and awarded categories.
@@
Figures 69 illustrate the factors influencing post-treatment perceptions according to the linear regression model with interaction terms (Eq. 2.2) for female and male participants separately. Post-treatment perceptions are shaped by different persistent factors. For female participants, pre-treatment perceptions of the manager’s gender-related trustworthiness, competence, and fairness significantly predict the corresponding post-treatment perceptions. In contrast, for male participants, pre-treatment perceptions of the manager’s type play a stronger role in shaping the post-treatment perceptions of all four dimensions. Demographic variables such as age and education generally show no significant effect on participants’ perceptions of their manager’s decision, except for male participants, where a postgraduate degree is associated with higher values of post-treatment ratings of managerial competence and fairness.

### v3

@@ Abstract
As AI becomes more embedded in workplaces, it is shifting from a tool for efficiency to an active force in organizational decision-making. Whether due to anthropomorphism or intentional design choices, people often assign human-like qualities, including gender, to AI systems. However, how AI managers are perceived in comparison to human managers and how gender influences these perceptions remains uncertain. To investigate this, we conducted randomized controlled trials (RCTs) where teams of three participants worked together under a randomly assigned manager. The manager was either a human or an AI and was presented as male, female, or gender-unspecified. The manager’s role was to select the best-performing team member for an additional award. Our findings reveal that while participants initially showed no strong preference based on manager type or gender, their perceptions changed notably after experiencing the award process. As expected, those who received awards rated their managers as more trustworthy, competent, and fair, and they were more willing to work with similar managers in the future. In contrast, those who were not selected viewed them less favorably. However, male managers, whether human or AI, were more positively received by awarded participants, whereas female managers, especially female AI managers, faced greater skepticism and negative judgments when they did not give awards. These results suggest that gender bias in leadership extends beyond human managers to include AI-driven decision-makers as well. As AI assumes more managerial responsibilities, understanding and addressing these biases will be crucial for designing fair and effective AI management systems.
@@ 1 Introduction
The increasing integration of artificial intelligence (AI) in workplaces is transforming its function from merely enhancing efficiency to actively shaping organizational decision-making, sparking discussions on the evolving dynamics between humans and machines [1]. AI algorithms are increasingly taking on roles traditionally performed by humans, such as assisting with decisions related to promotion and hiring [2, 3]. While human managers often retain final authority, some companies are experimenting with AI in leadership-adjacent roles, such as managing repetitive tasks or optimizing team operations [4]. The growing influence of AI raises questions about workplace dynamics, particularly how AI decision-makers are perceived relative to human managers. Issues such as fairness, biases, organizational justice [5], trust, morale, and willingness to collaborate within teams are becoming increasingly important as AI reshapes the future of leadership and management.
This study aims to examine how the combined effects of manager type (human or AI), manager gender, self-perceived contribution, and award outcome influence the perceptions of managers’ trustworthiness, competence, fairness, and participants’ willingness to work with a similar manager in future tasks. Trustworthiness relates to reliability and ethical standing; competence underpins a manager’s authority and efficacy; fairness is crucial for morale and acceptance; and willingness to work in future endeavors reflects the sustainability of the collaboration. These dimensions are essential for fostering positive and effective human-AI team interactions. By examining these dynamics, this study offers insights into how gendered presentations in AI managers impact human perceptions, ultimately contributing to a deeper understanding of effective AI integration in the workplace. The following section reviews prior research that informs our hypotheses.
2 Background
2.1 Gender bias in leadership evaluations
2.2 Anthropomorphism and gendering in AI systems
@@
Even when AI lacks explicit gender markers, individuals tend to subconsciously assign gender based on factors like language style, voice, and the perceived roles of AI. Research shows that minimal gender cues, such as vocal cues embedded in a machine, can evoke gender-based stereotypical responses [36]. A study on ChatGPT shows it is more likely to be perceived as male than female due to its competence-oriented attributes, such as providing information or summarizing text [25]. AI systems designed for caregiving or support, such as digital assistants like Siri and Alexa, tend to be feminized through their anthropomorphization [37], reflecting societal norms that associate caregiving and service roles with women [38]. The subconscious gendering of AI can influence perceptions of fairness and team dynamics, making it important to understand these biases to promote positive interactions in human-AI collaborations.
2.3 Fairness, self-perception and award outcomes
@@
2.4 Hypotheses
Together, the reviewed literature highlights how biases linked to manager type and gender can shape perceptions of managerial decision-making. Yet, the interplay between these factors, particularly in competitive contexts where only a subset of individuals receive recognition, remains underexplored. While gender bias is a central concern of this research, it is part of a broader investigation into how multiple interacting factors influence managerial perceptions. Building on prior literature and theoretical foundations, we propose the following hypotheses:
  • H1: Participants who receive an award will rate their manager as more trustworthy, competent, and fair, and express greater willingness to work with a similar manager in the future than participants who do not receive an award.
@@ 3 Materials and Methods
To explore these dynamics, we conducted online experiments where participants worked in teams under managerial scenarios varying by manager type and gender. The survey questions used in the pre-treatment and post-treatment surveys, along with screenshots from the experiment, are provided in the Supporting Information. This research complies with University College Dublin (UCD) Human Research Ethics Regulations and Human Research Ethics Committee (HREC) guidelines for research involving human participants. The research study protocol has been approved by UCD Human Research Ethics (HS-LR-24-10-Cui-Yasseri). Informed consent was obtained from all human participants prior to the commencement of the experiment.
@@ 3.2 Design and Measures
Figure 1 illustrates the timeline view of the experimental methodology. This study employed a mixed experimental design to examine how participants’ perceptions of trustworthiness, competence, fairness, and willingness to work with a manager were influenced by primary factors: manager type, manager gender, award outcome, and self-perceived contribution. Participants were randomly assigned to three-person groups, with each group placed in a different manager condition. In these conditions, participants viewed scenarios in which either a human or an AI manager made a selection decision. Within each manager type, the gender of the manager was also manipulated, with the manager identified as either female, male, or gender unspecified, to explore potential biases or differential expectations based on perceived gender.
@@
The dependent variables in this study were the participants’ perceptions of trustworthiness, competence, fairness of their assigned manager, and the participants’ willingness to collaborate with a similar manager in future tasks. Each of these perceptions was measured using a Likert-type scale from 0 to 10, with items such as “How fair do you think the manager was in making the award decision in the game?”. Take fairness as an example; we employed the following linear regression model to examine the factors influencing post-treatment fairness perceptions:
@@
3.3 Experimental task
Figure 2: Key partial views of the experiment interface. (A) Image of the individual task. (B) Manager assignment. (C) Team task. (D) Winner announcement.
The experiment consisted of an entry survey, two rounds of the “Find the Robber” task, and an exit survey. The task was adapted from a publicly available visual reasoning puzzle published on Bright Side [44]. The images and accompanying scenarios were modified to fit the experimental format and to simulate collaborative problem-solving and managerial evaluation. In the first round (individual round), each participant viewed an image accompanied by a short scenario and was asked to identify “the robber” from five cartoon suspects by adjusting a slider bar. They were required to submit the answer within a 2-minute limit. In the second round (team round), participants worked in teams of three and were asked to complete the same task collaboratively. A text-based chat interface allowed teammates to discuss their reasoning and jointly select a final team answer within a 4-minute limit. Figure. 2 shows key partial views of the experiment interface (see SI for complete screenshots and further information).
While the “Find the Robber” task is relatively simple and does not require advanced mathematics or logic, it engages key cognitive and social processes: perceptual discrimination, judgment under uncertainty (no obvious or guaranteed correct answer), time pressure, and social inference (how participants justify their reasoning and assess others’ input). Despite its simplicity compared with real-world workplace tasks, this task provides a controlled and interpretable model for studying reactions to award decisions in ambiguous, team-based settings where credit attribution is inherently uncertain.
@@ 3.4 Procedure
As shown in Fig. 1, the experiment followed a structured sequence of stages designed to simulate a team-based decision-making scenario involving managerial evaluation.
@@
Participant onboarding and pre-treatment survey. The experiment took place entirely online and was designed to simulate a decision-making scenario involving teamwork and managerial evaluation. Participants recruited via Prolific began by clicking a provided link to enter the Empirica platform (SI Fig. S1), where the experiment was hosted. Upon joining, participants first read an information sheet outlining the study’s goals and procedures, followed by providing their consent (SI Figs. S2-S3). After entering their Prolific ID, they proceed to the experiment (SI Figs. S4-S5). They then completed a brief survey collecting demographic information, such as age, gender, education level, and their initial perceptions of trustworthiness, competence, fairness, and willingness to work with both AI and human managers, as well as male and female managers. This initial survey helped establish baseline perspectives toward different manager types.
@@
Individual problem-solving task. After completing the preliminary survey, participants began the first round of the game. Each participant saw a countdown timer appeared at the top center of their screen, their randomly assigned ID, a short text description, a cartoon picture, a slider tool, and a submit button (SI Fig. S6). They were required to individually complete the visual intelligence task (“Find the Robber” ): read the on-screen description, examine the picture based on visual cues, and identify the robber from five options. Participants selected the suspect they believed to be the robber using the slider bar within a 2-minute time limit; otherwise, a default value was automatically recorded.
@@
Manager assignment. After completing the first round of the game, participants were informed they would join a team with two teammates for the next round (SI Figs S7–S10). Each team was randomly assigned a manager, either a human or an AI, who would evaluate the team’s performance. The manager’s gender was also manipulated within each manager type, with the manager identified as male, female, or gender unspecified. The manager would ultimately select the “best player” in the team to receive an extra cash award of £0.5.
@@
Team-based task and collaboration. Following the manager assignment, participants proceeded to the second round of the game. In the second round, participants collaborated in a team-based setting through an online chat interface to complete the same task as in the first round (SI Fig. S11). They could view their own and their teammates’ responses from the previous round on the slider bars, providing context for discussion. Through the chat, participants worked collectively to arrive at a final team response, simulating real-world team dynamics. All teams completed identical tasks, ensuring a consistent basis for comparing perceptions of the managers across groups.
@@
Manager decision and exit survey. After a waiting time of 20 seconds, participants received the announcements of the “best player” selected by the manager (SI Figs. S13–S14). Finally, participants completed an exit survey that included the study’s primary measures: perceptions of trustworthiness, competence, and fairness regarding the manager’s decision, as well as their willingness to work with a similar manager in the future.
Debriefing and payment. After completing the survey, participants were informed during the debriefing that no actual manager was involved in the experiment and that the selection was random. The entire experiment took approximately 12 minutes, providing an experience of teamwork, evaluation, and managerial decision-making. Participants were compensated at a rate of £9 per hour, prorated for the duration of the study. The chosen “best player” received an additional £0.5 bonus.
@@ 4.1 Comparison of pre- and post-treatment perceptions
Figure 3: Comparison of the mean perceived trustworthiness of different groups by manager type and gender. The upper row represents results for female participants, and the lower row represents results for male participants. (A) and (D) show the mean pre-treatment perceived trustworthiness. (B) and (E) depict the mean change in post-treatment perceived trustworthiness for awarded participants, while (C) and (F) illustrate the change for non-awarded participants. Error bars represent the standard error of the mean.
Figure 4: Comparison of the mean perceived competence of different groups by manager type and gender. The upper row represents results for female participants, and the lower row represents results for male participants. (A) and (D) show the mean pre-treatment perceived competence. (B) and (E) depict the mean change in post-treatment perceived competence for awarded participants, while (C) and (F) illustrate the change for non-awarded participants. Error bars represent the standard error of the mean.
Figure 5: Comparison of the mean perceived fairness of different groups by manager type and gender. The upper row represents results for female participants, and the lower row represents results for male participants. (A) and (D) show the mean pre-treatment perceived fairness. (B) and (E) depict the mean change in post-treatment perceived fairness for awarded participants, while (C) and (F) illustrate the change for non-awarded participants. Error bars represent the standard error of the mean.
Figure 6: Comparison of the mean expressed willingness to work with a similar manager in the future, by manager type and gender. The upper row represents results for female participants, and the lower row represents results for male participants. (A) and (D) show the mean pre-treatment perceived willingness. (B) and (E) depict the mean change in post-treatment perceived willingness for awarded participants, while (C) and (F) illustrate the change for non-awarded participants. Error bars represent the standard error of the mean.
In the pre-treatment survey, the mean of participants’ perceptions across manager types and genders was around 7, with slight variations (Fig. 36 A, D). Participants generally rated human managers more favorably than AI managers. Female participants gave slightly higher ratings to female managers, and male participants gave slightly higher ratings to male managers, though the overall gender differences were relatively small. The perceptions of managerial decisions show a weak positive correlation between human and AI managers but a strong positive correlation between female and male managers (See SI Fig. S11). This suggests that managerial perceptions in the pre-treatment phase may be more influenced by the manager’s type than by the manager’s gender.
@@
Conversely, not receiving an award (Fig. 36 C, F) generally reduced perceptions of managers, with the steepest declines observed for female-AI managers across both female and male participants, and a statistically significant effect among male participants (one-way ANOVA, $p<0.05$) on all dimensions. The dynamic shifts observed in evaluations before and after the treatment emphasize the influence of manager type, gender, and award outcomes.
@@ 4.2 Effects of demographics and pre-treatment perceptions
Figure 8: Coefficient estimates ($\beta$) for post-treatment perceived competence $Competence_{post}$ among female and male participants, based on pre-treatment perceptions, demographic factors, and interactions. Interaction terms capture the influence of manager type, manager gender, and participants’ self-perceived contributions on perceived competence. Statistical significance is denoted by $p<.05$. Error bars represent standard errors. Dashed lines visually separate pre-treatment factors, non-awarded AI categories, non-awarded human categories, and awarded categories.
@@
Figures 710 illustrate the factors influencing post-treatment perceptions according to the linear regression model with interaction terms (Eq. 3.2) for female and male participants separately. Post-treatment perceptions are shaped by different persistent factors. For female participants, pre-treatment perceptions of the manager’s gender-related trustworthiness, competence, and fairness significantly predict the corresponding post-treatment perceptions. In contrast, for male participants, pre-treatment perceptions of the manager’s type play a stronger role in shaping the post-treatment perceptions of all four dimensions. Demographic variables such as age and education generally show no significant effect on participants’ perceptions of their manager’s decision, except for male participants, where a postgraduate degree is associated with higher post-treatment ratings of managerial competence and fairness.